OpenStudy (anonymous):

The radius of a cylindrical gift box is (2x+3) in. The height of the gift box is twice the radius. What is the surface area of the cylinder? Write your answer as a polynomial form.

OpenStudy (raden):

SA = 2*pi*r^2 + 2*pi*r*h just put the values of r=2x+3 and h=2(2x+3)=4x+6

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

I'm still kinda confused, but thank you so much for you help.

OpenStudy (raden):

well the formula to find the surface area of cylinder is SA = 2*pi*r^2 + 2*pi*r*h known r=2x+3 and h=4x+6 so, SA = 2*pi*(2x+3)^2 + 2*pi*(2x+3)*(4x+6) = 2*pi*(4x^2+12x+9) + 2*pi*(8x^2 + 24x + 18) if u want simplify of it, it can be SA = 2*pi*(4x^2+12x+9) + 2*pi*(8x^2 + 24x + 18) =2*pi(4x^2+12x+9+8x^2 + 24x + 18) =2*pi(12x^2+36x+27) = (24x^2 + 72x + 54)pi inc